OpenStudy (anonymous):

factor the trinomial b² -14b +48

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, what are two numbers that go into 48 and add up to 14? @dontre

OpenStudy (igreen):

-14*

OpenStudy (anonymous):

8 and 6

OpenStudy (anonymous):

2

OpenStudy (freckles):

yes so you would have -8 and -6 since -8+(-6)=-14 and -8(-6)=48 so b^2-14b+48 will have a factorization that looks like (b+something)(b+another something) so replace the something and the other something with the numbers you found

OpenStudy (freckles):

that is pretend we have x^2-5x+6 x^2-5x+6 has factorization (x-3)(x-2) because -3(-2)=6 and -3+(-2)=-5

OpenStudy (anonymous):

(b-7) (b-24)?